jonahcam1347:

if u have a right triangle and the legs of the triangle r 5in, 7in, what is the let of he third angle

zunairah:

do you know the formulah

zunairah:

a^2 + b^2 = c^2

zunairah:

Your A and B would be 5 and 7

zunairah:

5 squared + 7 squared = Your missing value (C value)

jonahcam1347:

tru

zunairah:

so 25 + 49 = C

jonahcam1347:

yes

zunairah:

1 attachment
quackmafia:

i agree